# EveGuard: Defeating Vibration-based Side-Channel Eavesdropping with Audio Adversarial Perturbations

### v1

@@ 1 Introduction
We propose EveGuard, an innovative software-based defense mechanism to protect against privacy leakage from the loudspeaker-generated voice in SSEAs. As shown in Figure 1, EveGuard mitigates SSEA threats by introducing audio adversarial examples to the original audio signals prior to playback. EveGuard ensures that (i) the perturbed speech signals remain natural to human ears and microphones, and (ii) any attempt by SSEAs to capture and reconstruct the perturbed speech will produce content that is difficult to interpret, both for humans and automated speech recognition systems. Note that EveGuard cannot protect voice from a human speaker when SSEAs target eavesdropping on throat vibrations, a challenge also for state-of-the-art (SOTA) attacks [60].
@@ 2 Background and Related Work
2.1 Side-channel Speech Eavesdropping
@@
mmWave Radar-based SSEA. mmWave radars are commonly used for object ranging or imaging. Yet, slight object displacements can produce subtle phase changes in the reflected radar signals, enabling the reconstruction of audio signals from these changes [20, 21, 54, 62]. Although some hardware-enabled defense techniques [28, 71, 53, 56] exist, like jamming and IRS shield, these approaches typically provide only a limited area of defense and often fall short in effectiveness. For instance, IRS-based approaches [53, 56] generally require individuals targeted by SSEAs to procure and deploy extra hardware. The IRS device needs to be oriented toward the attacker and placed close (e.g., 1.2m [53]) to the attacker’s mmWave radar. Additionally, these defense mechanisms may unintentionally compromise the standard sensing capabilities of mmWave radars.
@@ 3.1 EveGuard Threat Model
We target the SOTA SSEA scenarios where the attacker has control of the side-channel sensors and can capture the speech-induced vibration on the loudspeakers. We assume that the attacker has strong capabilities to i). compromise the side-channel sensors and their sensing systems; ii). employ advanced ML and signal processing techniques to derive private information from the captured signals; iii). gain access to the victims’ speech samples or even create training datasets to facilitate the eavesdropping.
@@
For optical sensor-based SSEA, consistently with recently proposed attacks [49, 41], we assume the attacker can deploy their laser transmitter towards the reverberator without line-of-sight obstruction. The optical sensor captures the reflected laser, and the attacker accesses the ADC data, following similar assumptions as in mmWave-based SSEAs.
For the motion sensor-based SSEA, we assume that the attacker can trick the victim into installing a malicious app, which collects motion signals in the background and can even stream the data to the adversaries’ server.
@@ 4.2 Characterizing Human Hearing
We then investigate human auditory sensitivity across different frequency ranges to devise an undetectable defense. Our study is based on equal-loudness contour [2], a well-established model that delineates the sound pressure level (SPL) perceived by the human ear across the frequency bands. We convert SPL measurements from to into a normalized scale from to to visualize human insensitivity to different frequency bands. Figure 4 shows the hearing curve and the imperceptibility gain obtained through psychoacoustic experiments to describe the human ear’s sensitivity [2]. The hearing curve represents the amplitude required for a purely continuous tone of a certain frequency that humans can hear. Frequencies with high imperceptibility are harder to perceive by human. It shows that human ears are most sensitive to frequencies between 1.6 kHz and 4 kHz, with a marked insensitivity to frequencies below 500 Hz. Furthermore, we conducted an experiment to investigate the effects of low- and mid-frequency noise on human speech understanding. We asked 24 participants to listen to noisy audio containing either low-frequency or mid-frequency AWGN and then translate sentences. We calculated the WER based on their translations. From Fig 5(a), we observe that at SNRs above 6dB, low-frequency noise does not impact human understanding. However, as shown in Fig 5(b), for participants to accurately hear audio containing mid-frequency noise, the SNR must be over 16dB. Furthermore, SNR and defense success rate have an inverse relationship (see Sec. 6.7). Thus, to guarantee the sound quality to the human ear, EveGuard can mainly generate perturbations within the low-frequency bands (i.e., 500 Hz) while minimally affecting the mid- and high-frequency ranges (i.e., 500 Hz).
@@ 6.1 SSEA Setup and Implementation
i). Signal Processing (SP). We follow the steps discussed in Sec. 4.1 and [20] to pre-process the raw eavesdropping signals and improve reconstruction quality. Note that SP only applies to mmWave radar.
ii). Machine Learning (ML). To enhance the probability of successful eavesdropping, we further implement an ML-based speech enhancement model by following well-established cGAN models [21, 22].
iii). Speech Recognition (SR). The final step involves training dedicated speech recognition or audio classification models using the processed signals from SP and ML. It aims to extract explicit private information. The results of SP and ML in the evaluation section are derived through SR.
@@ 6.2 Defense Implementation
| Defense | ML-based SSEA | SP-based SSEA | ||||||
@@ 6.3 Evaluation Metrics
Mel-Cepstral Distortion (MCD) [33] quantifies the difference between the original speech and the attacker’s reconstruction. Reconstructed audio with an MCD below 8 is typically recognizable by speech recognition models [70].
@@
Perceptual Evaluation of Speech Quality (PESQ) [46] is a standardized metric to assess the speech quality. A score above 3.0 is required for good-quality voice communication.
@@
SSEA 1 Tx + 4 Rx MIMO (3 Txs + 4 Rxs) MCD WER DDR MCD WER DDR ML 13.4 13.9 SP 13.5 14.1
6.4 Results of Radar-based SSEA
@@
6.5 Results of Different SSEA Sensors
@@ 6.6 User Study for Speech Quality
We conduct a public survey to analyze the effect of EveGuard on voice quality. We recruited 24 volunteers aged 20-40 from different backgrounds, i.e., graduate students, educators, and industry professionals. The participants were requested to rate two criteria based on a Likert scale [37] from 1 to 5: (1) the intelligibility of the eavesdropper’s reconstructed audio and (2) perceptual quality of EveGuard’s perturbed audio. In order to avoid bias, participants are not informed that the audio is perturbed. The complete survey can be found in Appendix F. As shown in Figure 14(b), we found that 94.8 of participants could not discern any information from the SSEA-reconstructed audio, while 5.2 could vaguely hear a few words. Participants rated the naturalness of EveGuard’ perturbed audio highly, with an average score of 4.59, closely approaching the original audio’s perceptual quality score of 4.69, indicating that EveGuard preserves the audio quality.
| Original | Adversarial Training | Perturbation Subtraction | ||||||
| MCD | WER | DDR | MCD | WER | DDR | MCD | WER | DDR |
| 13.4 | 12.6 | 13.5 | ||||||
@@ 6.7 Ablation Study
Analysis of Eve-GAN. We evaluate the translation performance of Eve-GAN, which converts audio into the SSEA samples. To achieve this, we utilize the widely used Structural Similarity Index Measure (SSIM) [64] to quantify the spectrogram similarity between the generated and actual SSEA samples. We find that Eve-GAN achieves SSIM of 94.02 for mmWave and 96.51 for accelerometer, respectively. These high similarity rates show that Eve-GAN effectively produces samples that closely mirror real SSEA samples, thereby enabling the PGM to efficiently train on adversarial examples generated by Eve-GAN.
@@
| SSEA | Quantization | Audio Resampling | Frequency Filtering | ||||||
@@ 6.8 Runtime System Overhead of EveGuard
We evaluate the runtime latency of EveGuard on two platforms, including a workstation with NVIDIA RTX A6000 and a low-end desktop with RTX 2060. EveGuard converts audio with sampling rates of 16kHz and 48kHz into adversarial audio at 50ms granularity, respectively. Experimental results show that the high-end desktop experiences latency of 2.7ms and 6.5ms when the sampling rate is 16kHz and 48kHz, respectively. Additionally, the low-end desktop requires 4.6ms and 11.7ms to process 16kHz and 48kHz audio, respectively. Since the latency threshold for high-quality VoIP communications is set at 150ms [50], EveGuard is feasible for deployment in both cloud and on-device platforms. Furthermore, we investigate EveGuard performance based on the segment length of the audio fed to the PGM. From Figure 16, we see that MCD remains nearly constant for segment lengths above 25ms, indicating potential for improving latency overhead.
@@ 6.9 Robustness to Adaptive Attackers
We evaluate EveGuard against adaptive attackers who seek to further enhance ML-SSEA and SR-SSEA based on knowledge of EveGuard. We define an expert attacker who is aware that the loudspeaker is protected by EveGuard and knows the defense methodology. However, the attacker does not know the exact ML model of the Eve-GAN and PGM. We assume that the attacker has the eavesdropping results of a 100-second perturbation sample. The attacker then trains a substitute EveGuard with a different architecture (two more layers, different number of neurons) with his/her training data. We consider two types of strategies:
Adversarial Training. We aim to robustly train SSEA by allowing attackers to expand the training data using substitute PGM. Specifically, the attacker crafts his/her adversarial audio, performs SSEA to obtain the reconstructed audio, aggregates it to form a new dataset.
@@
Evaluation Results. We report experimental results in the baseline attack scenario of mmWave radar, and summarize the results for the adversarial training and perturbation removal in Table XII. We see that these attacks fail to mitigate the effects of our perturbations because the estimated perturbation used by the attacker has a different distribution from the actual perturbation. Even if the attacker has acquired some eavesdropping results for the actual perturbations, PGM enhances the diversity of perturbations, making it infeasible for attackers to train ML-SSEA and SR-SSEA that are robust to all perturbations. Table XIII shows the defense results against the transformation-based approaches. We observe that quantization and audio resampling introduce signal distortions in the eavesdropped audio, making ML-SSEA reconstruction worse. Frequency filtering can remove perturbations to some extent, but comes with the trade-off of degrading the original eavesdropping speech quality, resulting in high reconstruction errors.
@@ 7 Conclusion
Appendix A Additional Frequency Responses of mmWave Radar
@@ Appendix D ML Model for EveGuard
EveGuard operates under the black-box assumption, employing a surrogate model rather than the target model utilized by SSEA. Specifically, EveGuard implements audio enhancement via [29]. EveGuard adopts the LSTM-based speech recognition [55] and the audio classifier proposed in mmSpy [3] as surrogate ML models. Model architecture for few-shot Eve-GAN and PGM are described in Sec. 5.
@@ Appendix E Details of EveGuard Training
We provide comprehensive parameter settings for EveGuard training. Specifically, we assign and in Equation 2. In two-stage PGM training, we set , , in Equation 7. We utilize a total of surrogate models for ensemble training. We set to 16, which determines the signal power of LFAP.

### v2

@@ 1 Introduction
We propose EveGuard, an innovative software-driven defense mechanism to protect against privacy leakage from the loudspeaker-generated voice in SSEAs. As shown in Figure 1, EveGuard mitigates SSEA threats by introducing audio adversarial examples to the original audio signals prior to playback. EveGuard ensures that (i) the perturbed speech signals remain natural to human ears and microphones, and (ii) any attempt by SSEAs to capture and reconstruct the perturbed speech will produce content that is difficult to interpret, both for humans and automated speech recognition systems. Note that EveGuard cannot protect voice from a human speaker when SSEAs target eavesdropping on throat vibrations, a challenge also for state-of-the-art (SOTA) attacks [66].
@@ 2 Background and Related Work
2.1 Vibration-based Speech Eavesdropping
@@
mmWave Radar-based SSEA. mmWave radars are commonly used for object ranging or imaging. Yet, slight object displacements produce subtle phase changes in the reflected radar signals, enabling the reconstruction of audio signals from these changes [23, 24, 59, 68]. Although some hardware-enabled defense techniques [31, 79, 58, 61] exist, like jamming and IRS shield, these approaches typically provide only a limited area of defense and often fall short in effectiveness. For instance, IRS-based approaches [58, 61] generally require individuals targeted by SSEAs to procure and deploy extra hardware. The IRS device needs to be oriented toward the attacker and placed close (e.g., 1.2m [58]) to the attacker’s mmWave radar. Additionally, these defense mechanisms may unintentionally compromise the standard sensing capabilities of radars.
@@ 3.1 EveGuard Threat Model
The attacker’s goal is to eavesdrop on the audio content played by a loudspeaker. However, direct microphone deployment inside or outside the victim’s room is infeasible due to the following reasons: (a) hidden microphones inside the room can be easily detected by victims using spy detectors [44, 89], making covert placement ineffective and (b) if victims use loudspeakers in a soundproof room, external eavesdropping microphones cannot capture the audio. As a result, the attacker seeks to exploit SSEAs, leveraging side-channel sensors to capture speech-induced vibrations from the loudspeakers. Specifically, mmWave radar-based techniques [23, 59, 74] enable the attacker to penetrate soundproofing materials and eavesdrop on speech. Additionally, zero-permission motion sensors[63] can be exploited to infer private speech information from a smartphone. The attacker may also detect speech-induced vibrations by sensing reflected ambient light [46].
In this work, rather than limiting our focus to constrained SSEAs, we aim to develop defenses capable of mitigating even the worst-case scenarios, ensuring robustness against highly capable adversaries. We assume that the attacker has strong capabilities to i). compromise the side-channel sensors and their sensing systems; ii). employ advanced ML and signal processing techniques to derive private information from the captured signals; iii). gain access to the victims’ speech samples or even create training datasets to facilitate the eavesdropping.
@@
For optical sensor-based SSEA, consistent with recently proposed attacks [54, 46], we assume the attacker can deploy their laser transmitter towards the reverberator without line-of-sight obstruction. The optical sensor captures the reflected laser, and the attacker accesses the ADC data, following similar assumptions as in mmWave-based SSEAs.
For the motion sensor-based SSEA, we assume that the attacker can trick the victim into installing a malicious app, which collects motion signals in the background and can even stream the data to the adversaries’ server. Note that detecting the malicious app is a challenging problem [80], as attackers can effectively disguise it among legitimate apps, allowing them to bypass security measures.
@@ 4.2 Characterizing Human Hearing
We then investigate human auditory sensitivity across different frequency ranges to devise an undetectable defense. Our study is based on equal-loudness contour [2], a well-established model that delineates the sound pressure level (SPL) perceived by the human ear across the frequency bands. We convert SPL measurements from to into a normalized scale from to to visualize human insensitivity to different frequency bands. Figure 4 shows the hearing curve and the imperceptibility gain obtained through psychoacoustic experiments to describe the human ear’s sensitivity [2]. The hearing curve represents the amplitude required for a purely continuous tone of a certain frequency that humans can hear. Frequencies with high imperceptibility are harder to perceive by human. It shows that human ears are most sensitive to frequencies between 1.6 kHz and 4 kHz, with a marked insensitivity to frequencies below 500 Hz. Furthermore, we conducted an experiment to investigate the effects of low- and mid-frequency noise on human speech understanding. We asked 24 participants to listen to noisy audio containing either low-frequency or mid-frequency AWGN and then translate sentences. Please refer to Sec. 6.6 for a detailed discussion of the ethical considerations in our study. We calculated the WER based on their translations. From Fig 5(a), we observe that at SNRs above 6dB, low-frequency noise does not impact human understanding. However, as shown in Fig 5(b), for participants to accurately hear audio containing mid-frequency noise, the SNR must be over 16dB. Furthermore, SNR and defense success rate have an inverse relationship (see Sec. 6.7). Thus, to guarantee the sound quality to the human ear, EveGuard can mainly generate perturbations within the low-frequency bands (i.e., 500 Hz) while minimally affecting the mid- and high-frequency ranges (i.e., 500 Hz).
@@ 6.1 SSEA Setup and Implementation
- 1.
- 2.
3.
Speech Recognition (SR). The final step involves training dedicated speech recognition or audio classification models using the processed signals from SP and ML. It aims to extract explicit private information. The results of SP and ML in the evaluation are derived through SR.
@@ 6.2 Defense Implementation
| Defense | ML-SSEA | SP-SSEA | ||||||
@@ 6.3 Evaluation Metrics
- 1.
2.
@@
4.
Perceptual Evaluation of Speech Quality (PESQ) [51] is a standardized speech quality metric. A score above 3.0 is required for good-quality voice communication.
@@
SSEA 1 Tx + 4 Rxs MIMO (3 Txs + 4 Rxs) MCD WER DDR MCD WER DDR ML 13.4 13.9 SP 13.5 14.1
6.4 Results of mmWave Radar-based SSEA
@@
6.5 Results of Different Sensor-based SSEAs
@@ 6.6 User Study for Speech Quality
In this study, we recruited 24 volunteers to evaluate the naturalness of audio samples. Specifically, we conducted an online survey in which participants listened to 30 pairs of audio samples, comprising a balanced mix of clean speech, perturbation-added speech, and speech eavesdropped from these two types, totaling 120 samples. To mitigate bias, the sample sequence was randomized, and the study’s objective was concealed. Participants were also encouraged to remain engaged throughout the evaluation. We received an Institutional Review Board (IRB) exemption from our institute (IRB 812263), confirming compliance with ethical guidelines. We collected limited demographic information, including participants’ age range (20–40) and professional backgrounds: 58.3 graduate students, 33.3 industry professionals, and 8.3 educators. No personally identifiable information (PII) was collected.
The participants were requested to rate two criteria based on a Likert scale [41] from 1 to 5: (1) the intelligibility of the eavesdropper’s reconstructed audio and (2) perceptual quality of EveGuard’s perturbed audio. In order to avoid bias, participants are not informed that the audio is perturbed. The complete survey can be found in Appendix F. As shown in Figure 14(b), we found that 94.8 of participants could not discern any information from the SSEA-reconstructed audio, while 5.2 could vaguely hear a few words. Participants rated the naturalness of EveGuard’ perturbed audio highly, with an average score of 4.59, closely approaching the original audio’s perceptual quality score of 4.69, indicating that EveGuard preserves the audio quality.
| of Samples | Original | AT | PR | ||||||
| MCD | WER | DDR | MCD | WER | DDR | MCD | WER | DDR | |
| 100 secs | 13.4 | 12.6 | 13.5 | ||||||
| 30 mins | 13.4 | 11.8 | 12.8 | ||||||
@@ 6.7 Ablation Study
Analysis of Eve-GAN. We evaluate the translation performance of Eve-GAN, which converts audio into the SSEA samples. To achieve this, we utilize the widely used Structural Similarity Index Measure (SSIM) [71] to quantify the spectrogram similarity between the generated and actual SSEA samples. The test dataset contains 5,600 audio samples. We find that Eve-GAN achieves SSIM of 94.02 for mmWave and 96.51 for accelerometer, respectively. The standard deviation for radar and accelerometers is 1.54 and 2.27. These high similarity rates show that Eve-GAN effectively produces samples that closely mirror real SSEA samples, thereby enabling the PGM to efficiently train on adversarial examples generated by Eve-GAN.
@@
| SSEA | QT | AR | FF | ||||||
@@ 6.8 Runtime System Overhead of EveGuard
We evaluated the run-time latency of EveGuard on two hardware platforms, including a workstation with NVIDIA RTX A6000 and a low-end desktop with RTX 2060. EveGuard converts audio with sampling rates of 16kHz and 48kHz into adversarial audio at 50ms granularity, respectively. Experimental results show that the high-end desktop experiences latency of 2.7ms and 6.5ms when the sampling rate is 16kHz and 48kHz, respectively. Additionally, the low-end desktop requires 4.6ms and 11.7ms to process 16kHz and 48kHz audio, respectively. Since the latency threshold for high-quality VoIP communications is set at 150ms [55], EveGuard is feasible for deployment in both cloud and on-device platforms. Furthermore, we investigate EveGuard performance based on the segment length of the audio fed to the PGM. From Figure 16, we see that MCD remains nearly constant for segment lengths above 25ms, indicating potential for improving latency overhead. As a further improvement, we will consider offloading computation from the loudspeaker to a cloud for collaborative defense.
@@ 6.9 Robustness to Adaptive Attackers
We evaluate EveGuard against adaptive attackers who seek to further enhance ML-SSEA and SR-SSEA based on knowledge of EveGuard. We define an expert attacker who is aware that the loudspeaker is protected by EveGuard and knows the defense methodology. However, the attacker does not know the exact ML model of the Eve-GAN and PGM. We assume that the attacker has access to multiple eavesdropping results from original-perturbed sample pairs. To observe the impact of sample quantity on defense performance, we consider two scenarios where the length of sample pairs corresponds to 100 seconds and 30 minutes. The attacker then trains a substitute EveGuard with a different architecture (two more layers, different number of neurons) with his/her training data. Additionally, we explore an attack strategy in which the attacker may apply transformation operations to mitigate the perturbation effects. Overall, we consider three types of strategies:
Adversarial Training. We aim to robustly train SSEA by allowing attackers to expand the training data using substitute PGM. Specifically, the attacker crafts his/her adversarial audio, performs SSEA to obtain the reconstructed audio, and aggregates it to form a new dataset.
@@
Evaluation Results. We report experimental results in the baseline attack scenario of mmWave radar, and summarize the results for the adversarial training and perturbation removal in Table XII. We see that these attacks fail to mitigate the effects of our perturbations because the estimated perturbation used by the attacker has a different distribution from the actual perturbation. Even if the attacker has acquired some eavesdropping results for the actual perturbations, PGM enhances the diversity of perturbations, making it infeasible for attackers to train ML-SSEA and SR-SSEA that are robust to all perturbations. Table XII presents a performance comparison between two scenarios in which the attacker has access to original-perturbed sample pairs of different lengths. The results indicate a slight improvement in the attacker’s eavesdropping when given access to a large volume of EveGuard’s samples; however, the overall eavesdropping capability remains highly limited. This is because the attacker lacks full knowledge of EveGuard, and the PGM-generated perturbations are highly diverse with random triggers, making adaptive attacks challenging. This observation aligns with the analysis in [10, 3].
Table XIII shows the defense performance against speech transformation-based approaches. We observe that quantization and audio resampling introduce signal distortions in the eavesdropped audio, making ML-SSEA reconstruction worse. Frequency filtering can remove perturbations to some extent, but comes with the trade-off of degrading the original eavesdropping speech quality, resulting in high reconstruction errors.
7 Discussion
Electromagnetic (EM) Side-Channel Eavesdropping. Several studies [65, 69, 14, 81, 40] have explored audio eavesdropping via EM side channels, which can be categorized into two strategies: (a) passive EM sensing and (b) active EM sensing attacks. Specifically, passive EM sensing [14, 40] exploits unintended EM emissions generated when playing audio. Since EM leakage is inherently weak and difficult to detect remotely, passive attacks require close proximity (around 50 cm) to the target. Active EM sensing attacks [65, 69, 81] involve embedding a retroreflector inside the target device, which reflects the attacker’s RF signal back to a receiver, enhancing the leakage for analysis. However, this approach relies on stronger assumptions, as it requires prior physical access to the target.
Overall, EM side-channel eavesdropping poses weaker security risks compared to vibration-based SSEAs. As part of our future work, we aim to extend EveGuard to defend against EM side-channel eavesdropping by designing universal perturbations. This approach leverages the frequency response differences between EM side channels and traditional microphones.
Eavesdropping on Human Speech. Our focus is on protecting audio played through loudspeakers. Prior research [66] has experimentally demonstrated that eavesdropping on live human speech is unlikely in real-world scenarios. However, reactive defense mechanisms, such as playing audio perturbations upon detecting speech, could be explored as a potential countermeasure, which we leave for future work.
@@ 8 Conclusion
Acknowledgement
We thank the anonymous reviewers and our shepherd for their invaluable comments and feedback that helped us improve our manuscript. This work is partially supported by the U.S. Army/Department of Defense award number W911NF2020267 and Google Ph.D. Fellowship.
Appendix A Additional Frequency Responses
@@ Appendix D ML Model for EveGuard
EveGuard operates under the black-box assumption, employing a surrogate model rather than the target model utilized by SSEA. Specifically, EveGuard implements audio enhancement via [32]. EveGuard adopts the LSTM-based speech recognition [60] and the audio classifier proposed in mmSpy [4] as surrogate ML models.
Eve-GAN is composed of five key components: a content encoder, a domain encoder, a bottleneck extractor, a generator, and a discriminator. The content encoder maps human speech into a latent content code, while the domain encoder extracts SSEA-specific features. The bottleneck extractor refines these representations before feeding them into the generator, which reconstructs the audio using AdaIN residual blocks and upscale Conv1d layers. The discriminator, based on a multi-period architecture [37], evaluates the realism of the generated audio across multiple frequency scales to enhance robustness against SSEA-based attacks.
The PGM architecture consists of three main modules. The FIR generator learns the audio perturbation distribution in the frequency domain. The LFAP generator induces low-frequency perturbations to exploit the unique frequency response characteristics of sensors, preventing attackers from accurately restoring speech. Finally, the discriminator distinguishes between real and generated audio, ensuring that perturbations remain imperceptible while making it difficult for adaptive attackers to learn the patterns of EveGuard.
@@ Appendix E Details of EveGuard Training
We provide comprehensive parameter settings for EveGuard training. Specifically, we assign and in Eq. 2. In two-stage PGM training, we set , , in Eq. 7. We utilize a total of surrogate models for ensemble training. We set to 16, which determines the signal power of LFAP.
@@ Appendix G Details of Speech Transformation
Appendix H Meta-Review
The following meta-review was prepared by the program committee for the 2025 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
H.1 Summary
This paper introduces EveGuard, a software-based defense mechanism designed to protect against vibration-based side-channel eavesdropping attacks. These attacks exploit sensors such as radars, accelerometers, and optical sensors. The primary contributions include the Perturbation Generator Module (PGM), which generates adversarial perturbations to disrupt side-channel attacks without significantly degrading audio quality for human listeners, and Eve-GAN, a generative adversarial network trained with minimal data to simulate various attack scenarios, thereby enhancing generalizability and practicality.
H.2 Scientific Contributions
Creates a New Tool to Enable Future Science
Provides a Valuable Step Forward in an Established Field
H.3 Reasons for Acceptance
1.
Creates a New Tool to Enable Future Science: EveGuard presents a software-based approach for defending against vibration-induced side-channel attacks. The methodology, particularly the use of Eve-GAN and adversarial perturbations, is robust and convincingly validated through comprehensive experimentation.
2.
Provides a Valuable Step Forward in an Established Field: EveGuard effectively addresses a recognized security threat, improving the flexibility and applicability of existing defense solutions. The proposed PGM and Eve-GAN end-to-end design improves both defense effectiveness and usability.